# Effort-aware Fairness: Incorporating a Philosophy-informed, Human-centered Notion of Effort into Algorithmic Fairness Metrics

### v1

@@ Abstract
Although popularized AI fairness metrics, e.g., demographic parity, have uncovered bias in AI-assisted decision-making outcomes, they do not consider how much effort one has spent to get to where one is today in the input feature space. However, the notion of effort is important in how Philosophy and humans understand fairness. We propose a philosophy-informed way to conceptualize and evaluate Effort-aware Fairness (EaF) based on the concept of Force, or temporal trajectory of predictive features coupled with inertia. In addition to our theoretical formulation of EaF metrics, our empirical contributions include: 1/ a pre-registered human subjects experiment, which demonstrates that for both stages of the (individual) fairness evaluation process, people consider the temporal trajectory of a predictive feature more than its aggregate value; 2/ pipelines to compute Effort-aware Individual/Group Fairness in the criminal justice and personal finance contexts. Our work may enable AI model auditors to uncover and potentially correct unfair decisions against individuals who spent significant efforts to improve but are still stuck with systemic/early-life disadvantages outside their control.
@@ Introduction
AI has assisted humans in making high-stakes decisions such as recidivism risk assessment (e.g., Bao et al. 2021) and loan approval (e.g., Mayer, Strich, and Fiedler 2020), leading to potentially increased efficiency and consistency. However, the growing deployment of AI in critical decision-making contexts has also raised serious concerns about algorithmic bias, which can disproportionately impact individuals and communities. In response, the AI literature has deeply explored two notions of fairness: group fairness (ensuring that a quantity of interest, such as accuracy, should be equalized across demographic groups like race/sex) (e.g., Pedreshi, Ruggieri, and Turini 2008) and individual fairness (ensuring that similar individuals get similar outcomes) (e.g., Dwork et al. 2012).
@@
Typical applications of both group and individual fairness treat similarly situated individuals as deserving similar treatment despite potential differences in demographic features (e.g., race, sex, and age…). However, similarly situated individuals, e.g., those with the same income or arrest history, are often similarly situated for very different reasons. This leads to a hypothetical “identical CVs” problem: two people from very different backgrounds may have the same CV—and, therefore, are treated equivalently per group or individual fairness—but one of them may have had to work a lot harder to achieve that. The notion that effort is an important factor in fairness has been well studied in philosophy (e.g., the free rider problem Hardin and Cullity 2003) but has been largely, though not completely (see related work below), absent from the AI fairness literature. One key ingredient of efforts is the temporal dimension: How an individual has changed over time. Liu et al. (2018) show that in an optimization problem, imposing a static fairness constraint, which often ignores the temporal dimension, may harm long-term improvement.
@@
We first conceptualize a new measure of Effort-aware Fairness (EaF), grounded in philosophical literature (De Tracy 1803; Massin 2017; Maine de Biran 2002), based on two key ideas: 1/ acceleration, or more broadly the temporal trajectory of how an individual’s task-relevant feature has changed over time; 2/ inertia or how much historical disadvantage each individual may have experienced outside their controls. We then conduct a human subjects experiment to validate whether our trajectory-based formulation of effort is more relevant to laypeople’s fairness perception than traditional formulations based solely on the feature’s aggregate value (Romano, Bates, and Candes 2020; Ni et al. 2024). Finally, we demonstrate how our EaF metric can be computed to audit AI models on two datasets, CLUE (criminal justice) and SHED (personal finance).  [^1] Our supplementary materials (such as record-linked SHED data, human subjects experiment data, Python/Jupyter Notebook code, survey print-out) are uploaded to an anonymized Google Drive folder: https://drive.google.com/drive/folders/1RsSE1dUHDkn41iKCkIhY5QRdAVKVcrtB?usp=sharing
@@
In the technical CS/AI fairness literature, there exist two main approaches to mathematically formulate “fairness” based on the outcome of human/AI-assisted decision-making systems: group fairness and individual fairness. Group fairness, or group parity, is achieved when a statistical metric of interest (such as positive outcome rate or false positive rate) is equalized across different groups with respect to a sensitive feature, e.g. race or sex (Pedreshi, Ruggieri, and Turini 2008; Barocas, Hardt, and Narayanan 2019). Individual fairness, introduced formally by Dwork et al. (2012), is based on the intuition that similar individuals should get similar outcomes. Formally, for a pair of individuals, given an input-space distance metric to measure how differently situated they are and an output-space distance metric to measure how different the distributions of their possible outcomes are, individual fairness requires (as a constraint in an optimization problem) that their output-space distance should be upper-bounded by their input-space distance. The more similarly situated two persons are (e.g., same demographics and criminal histories), the more likely that they should receive similar outcomes (e.g., risk scores).  [^2] We discuss the legal relevance of group fairness and individual fairness under U.S. law in Appendix A.
@@
Recent AI fairness works attempt to loop some notion of “effort” into fairness metrics by equalizing “effort” across groups. They conceptualize “effort” as the relationship between changes in mutable input features (e.g., prior arrests) and the probability predicted by a model of receiving a favorable outcome (e.g., granting bail). Approaches differ in terms of whether they take outcome parity as a requirement and attempt to equalize “effort” across groups (Heidari, Nanda, and Gummadi 2019; Guldogan et al. 2023) or whether they take equalized “effort” as a requirement and attempt to equalize outcome parity (Gupta et al. 2019; Huang et al. 2020; Von Kügelgen et al. 2022). These dual styles of approach are closely related to work on algorithmic recourse (Ustun, Spangher, and Liu 2019), in that they ask how much “effort” an individual would have to put in in order to achieve a favorable outcome. Both approaches can be loosely considered to be in the space of counterfactual “effort” and in contrast to our approach—we are interested in how much “effort” an individual has already put in and use that as the value against which we seek parity.
Formally, existing approaches focus on counterfactual changes in input features ($\Delta\mathbf{x}$), which may lead to changes in the probability of receiving a favorable output label, e.g. from $f(\mathbf{x})<\gamma$ to $f(\mathbf{x}+\Delta\mathbf{x})\geq\gamma$ where $\gamma$ is a probability threshold. On the one hand, Heidari, Nanda, and Gummadi (2019) and Guldogan et al. (2023) set an upper threshold $\delta$ for the norm function $\mu$ of “efforts”, i.e. $\mu(\Delta\mathbf{x})\leq\delta$, and seeks to equalize across groups the maximized favorable output rate. On the other hand, Gupta et al. (2019), Huang et al. (2020), and Von Kügelgen et al. (2022) set a lower threshold for the desirable output rate and seeks to equalize across groups the minimized norm of “efforts”. However, these works do not use actual changes over time, but they compare an individual with all others within the same group at only one time step to estimate $\Delta\mathbf{x}$ and $f(\mathbf{x}+\Delta\mathbf{x})$.  [^3] Shortcomings of the counterfactual effort fairness literature and how our formulation overcomes them are in Appendix B.
@@
As reviewed by Massin (2017), the philosophy literature has used four main approaches to characterize effort, of which two are directly related to effort itself (rather than the perception of effort): 1/ Force-based account (how much force would it take to “move” an object; introduced by De Tracy (1803) and later formalized/advocated for by, i.a., Maine de Biran (2002)) and 2/ Resource- or Energy-based account (spending effort towards a goal means allocating part of a finite amount of resources, or energy tank, up until the energy tank is depleted; introduced by Arai (1912) and later formalized/advocated for by, i.a., Gendolla, Wright, and Richter (2012)). The Force-based account addresses several problems unresolved by the Energy-based account, such as capturing the concept of “resistance” to effort, distinguishing between “fatigue” and “effort”, and explaining why effort is unpleasant but praiseworthy. Massin (2017) informedly concludes that the Force-based account is “more fundamental”.
The three main Philosophy arguments by Massin (2017) in support of the force-based account are highly applicable to our AI fairness contexts: 1/ As “only force-based accounts properly capture the idea of a resistance to our effort,” these accounts can capture issues such as systemic discrimination that resists individual-level progress (Bohren, Hull, and Imas 2022). 2/ As “force-based accounts, by contrast [to energy/resource-based accounts], stay clear of any substantive commitment about the relationship between fatigue and effort,” it is compatible with our AI-assisted decision-making contexts where a decision subject with better effort (e.g., fewer arrests or higher income after each year) does not necessarily suffer from more fatigue. 3/ As “force-based accounts neatly explain the apparently paradoxical facts that we seem to praise and enjoy efforts partly in virtue of their unpleasantness,” it is compatible with the later introduced inertia and acceleration terms in our force-based formulation of effort because having grown up in childhood poverty but still managing to work harder for better income (or desistance) might be unpleasant but indicating of virtue and thus praiseworthy (McNeill and Farrall 2013). This force-based conceptualization of effort is later expanded to conceptualize the “feeling of effort” by Bermúdez and Massin (2023): “Among the possible effort-first accounts, a force-based approach seems most promising, partly because it allows for a direct explanation of why the feeling of effort involves the experience of countering a resistance.”
@@
To estimate “effort” as an individual-level metric, we formulate the “more fundamental” Force-based account of “effort” from philosophy (Massin 2017) by using Newton’s Second Law of Motion from Physics to develop an Effort-as-Force metric. Building on this idea, we formulate new Effort-aware Fairness metrics to evaluate on real-world datasets.
@@
[…]
@@
Our first quantity to model is inertia $m$. We conceptualize $m$ as the societal “holding-back” effect and/or historical disadvantage against an individual beyond their control/agency. One example metric for inertia is hereditary disability, i.e., a disability passed down to a person by their parents. However, this feature might be hard to obtain (due to health privacy concerns) and hard to assign a numeric value. Another example metric for inertia is the socio-economic background of a person’s family during their childhood, e.g., whether back when the person was still a child, their family had to live below the federal poverty line, which is beyond the child’s control. In our criminal justice and personal finance contexts, childhood poverty was found to be associated with significantly higher “hazard rate of being convicted of violent criminality” (Sariaslan et al. 2014) and to have “quantitatively large detrimental effects” on adult earnings (Duncan, Ziol-Guest, and Kalil 2010).
@@
Next, acceleration $\mathbf{a}$ can be modelled as the second-order time derivative of task-relevant input features (e.g., number of past arrests in criminal justice, or income in personal finance). However, at first sight, there seems to be a problem with second-order derivative. Suppose over a three-year period with equidistant time intervals ($\Delta t=\text{one year}=1$), a person’s annual income increases from $x_{0}=60k$ USD (first year) to $x_{1}=90k$ USD (second year) to $x_{2}=100k$ USD (third year). Our intuition would stipulate that they demonstrate positive effort overall as their annual income kept increasing, despite with slowing velocity ($v_{0}=90k-60k=30k$ USD/year followed by $v_{1}=10k$ USD/year). However, the slowing speed would result in a negative second-order derivative of $a_{0}=10k-30k=-20k$ (USD/year2), counterintuitively implying negative efforts. Does this counterintuitive result mean the Force-based approach for Effort is wrong? No! The underlying issue is that we should have modeled the temporal input feature (income) in a cumulative view: $X_{0}=60k$ USD, $X_{1}=60k+90k=150k$ USD, and $X_{2}=150k+100k=250k$ USD. In this cumulative formulation of income, velocity terms are $V_{0}=90k$ (USD/year) and $V_{1}=100k$ (USD/year); acceleration is $A_{0}=10k$ (USD/year2), intuitively implying positive efforts. We use lowercase letters ($x,v,a$) to signal a non-cumulative view and uppercase letters ($X,V,A$) to signal a cumulative view. This example highlights the importance of transforming a non-cumulative feature into a cumulative feature before computing its acceleration to quantify effort.
@@
We justify formally our choice of a cumulative formulation of the effort-related feature to compute acceleration. A non-cumulative version of a feature can be obtained by differentiating (taking the derivative of) the cumulative version of that same feature. Therefore, acceleration (second-order derivative, i.e., derivative of the first-order derivative) in the cumulative version corresponds to velocity (first-order derivative) in the non-cumulative version, e.g., $A_{0}=v_{1}$. If we took the second-order derivative in the non-cumulative version, it would correspond to the third-order derivative in the cumulative version, which is theoretically at odds with force and hard to compute in practice. Since effort and AI-assisted decision making are social phenomena, we explain the social science relevance of our force-based formulation of effort in Appendix C.
@@
To develop an Effort-aware Individual Fairness (EaIF) metric, we stipulate that similar individuals who spent similar efforts should get similar outcomes. We propose a pairwise individual similarity function as a weighted combination between the aggregate task-relevant feature and effort $E$.
We use Definition 2.1 (Lipschizt mapping) by Dwork et al. (2012) as the individual fairness condition. Let $\mathbf{x}^{i}$ and $\mathbf{x}^{j}$ denote task-relevant feature vectors for a pair of individuals. A risk predictor $M$ maps individual $\mathbf{x}^{i}$ to a risk score, i.e., $M\mathbf{x}^{i}$. We define an effort function $E(\cdot)$ that maps individual $\mathbf{x}^{i}$ to an effort score $E(\mathbf{x}^{i})$ and an aggregate (summing) function $S(\cdot)$ (e.g., to sum all non-cumulative values of income or past arrests $S(\mathbf{x}^{i})$ of individual $\mathbf{x}^{i}$). We also define an input-space individual distance function $d(\cdot,\cdot)$ over individuals’ feature representations and an output-space distance function $D(\cdot,\cdot)$ over their predicted risk scores. The predictor $M$ achieves perfect effort-aware individual fairness if for every pair of ($\mathbf{x}^{i},\mathbf{x}^{j}$), this constraint is satisfied:
@@
[…]
@@
where $[e_{0},e_{0}+\epsilon]$ is a small range (bin) of effort values, which can be computed from inertia $m$ and acceleration $\mathbf{A}$ (from cumulative formulation of effort-related feature). The terms $g_{i}$ and $g_{j}$ are demographic groups (e.g., White and Black) and $\mathcal{G}$ denotes the set of all such groups.
@@
For the next two empirical sections (human subjects experiment and metric demonstration pipelines), we use two real-world datasets that are relevant to two different AI-assisted decision-making context and have a temporal dimension. We also build simple predictive models to obtain AI predictions, on which Effort-aware Fairness can be evaluated.
@@
We draw defendants from a real dataset, Client Legal Utility Engine (CLUE), with over $4.4$ million District Court and $700,000$ Circuit Court cases between $2012$ and $2020$ scraped by the Maryland Volunteer Lawyers Service (MVLS) from the Maryland Judiciary Case Search. [^5] https://casesearch.courts.state.md.us/casesearch/ CLUE contains criminal history and demographic features, e.g. race, sex, and age (at the time of arrest).
@@
For every defendant who had an arrest between $2016$ and $2018$, we develop simple machine learning models (Random Forest classifier, whose predictions we use for the human subjects experiment, and three other models, LightGBM, Logistic Regression, and Decision Tree, whose predictions we use for the metric demonstration) with 5-fold cross-validation to predict their (violent) recidivism risk (on a scale of 00 to $1$) within one year current their recent arrest. [^6] We provide more details about the predictive models in Appendix D.
@@
We use the “Survey of Household Economics and Decisionmaking” (SHED) published by the Federal Reserve since 2013.777www.federalreserve.gov/consumerscommunities/shed.htm To obtain the time dimension, we perform record linkage across four years ($2019$ to $2022$). After selecting only the households that have data available across all four years, we obtain the final set of $704$ households. Our effort-related feature is the annual range of household income (I40). We turn this range-based feature into a number by taking the lower bound. Our outcome feature of interest is the respondent’s frequency of unpaid credit card balance (C4A), which can be of real-world predictive interest for financial institutions to process credit/loan applications.
@@
We must first decide on an appropriate population. Several major works in the AI fairness literature have used laypeople’s perception of fairness to decide on an appropriate fairness metric for a specific task, or even to construct new fairness metrics. Saxena et al. (2019) investigate which of three fairness definitions (individual fairness, fairness in learning, calibrated fairness) laypeople (MTurk crowdworkers) perceive to be the fairest in the context of loan decisions. Harrison et al. (2020) conduct survey-based experiments where laypeople (MTurk crowdworkers) were asked to compare two models representing trade-offs between fairness-related properties. Wang et al. (2019) leverage metric learning to advance individual fairness by incorporating laypeople’s (Prolific crowdworkers’) fairness annotations in the COMPAS recidivism prediction task. We follow these precedents of informing our fairness metrics via laypeople’s fairness perception, based on two rationales. In principle, laypeople’s fairness judgment is important for policy. Awad et al. (2020) argue that policymakers should be aware of and prepare for societal pushback of potentially beneficial but high-stakes decision-making AI tools. Liu, Du, and Li (2021) maintain that the discrepancy between a human-AI legal responsibility allocation framework from laws and public’s expectations may hinder the adoption of and trust in AI at large. In practice, to potentially achieve statistical significance, a human subjects experiment typically needs hundreds of participants. For example, the mean sample size for online human subjects experiments at CHI, a leading Human-Computer Interaction conference, was 224 participants (Caine 2016). It is both expensive and logistically impractical to recruit such a large sample of technical CS/philosophy domain experts.

### v3

@@ Abstract
Although popularized AI fairness metrics, e.g., demographic parity, have uncovered bias in AI-assisted decision-making outcomes, they do not consider how much effort one has spent to get to where one is today in the input feature space. However, the notion of effort is important in how Philosophy and humans understand fairness. We propose a philosophy-informed approach to conceptualize and evaluate Effort-aware Fairness (EaF), grounded in the concept of Force, which represents the temporal trajectory of predictive features coupled with inertia. Besides theoretical formulation, our empirical contributions include: (1) a pre-registered human subjects experiment, which shows that for both stages of the (individual) fairness evaluation process, people consider the temporal trajectory of a predictive feature more than its aggregate value; (2) pipelines to compute Effort-aware Individual/Group Fairness in the criminal justice and personal finance contexts. Our work may enable AI model auditors to uncover and potentially correct unfair decisions against individuals who have spent significant efforts to improve but are still stuck with systemic disadvantages outside their control.
@@ Introduction
AI has assisted humans in making high-stakes decisions such as recidivism risk assessment (e.g., Bao et al. 2021) and loan approval (e.g., Mayer, Strich, and Fiedler 2020), leading to potentially increased efficiency. [^1] By 2023, $42\%$ of enterprises have actively deployed AI, while another $40\%$ are experimenting with the technology. Among those using or exploring AI, $59\%$ have accelerated their investments and implementation Efforts over the past two years (IBM 2024). However, the growing deployment of AI in critical decision-making contexts has also raised serious concerns about algorithmic bias, which can disproportionately impact individuals and communities. In response, the AI literature has deeply explored two notions of fairness: group fairness (ensuring that a quantity of interest, such as accuracy, is equalized across demographic groups like race/sex) (e.g., Pedreshi, Ruggieri, and Turini 2008) and individual fairness (ensuring that similar individuals get similar outcomes) (e.g., Dwork et al. 2012).
@@
Typical applications of group and individual fairness treat similarly situated individuals as deserving similar treatment despite potential differences in demographic features (e.g., race, sex, and age). However, similarly situated individuals, e.g., those with the same income or arrest history, are often similarly situated for very different reasons. This leads to a hypothetical “identical CVs” problem: two people from very different backgrounds may have the same CV—and, therefore, are treated equivalently per group or individual fairness—but one of them may have had to work a lot harder to achieve that. The notion that Effort is an essential factor in fairness has been well studied in philosophy (e.g., the free rider problem Hardin and Cullity 2003) but has been largely, though not completely (see related work below), absent from the AI fairness literature. One key ingredient of Efforts is the temporal dimension: How an individual has changed over time. Liu et al. (2018) show that in an optimization problem, imposing a static fairness constraint, which often ignores the temporal dimension, may harm long-term improvement.
@@
We first conceptualize a new measure of Effort-aware Fairness (EaF), grounded in philosophical literature (De Tracy 1803; Massin 2017; Maine de Biran 2002), based on two key ideas: (1) acceleration, or more broadly the temporal trajectory of how an individual’s task-relevant feature has changed over time; (2) inertia, or how much historical disadvantage each individual may have experienced outside their controls. We then conduct a human subjects experiment to validate whether our trajectory-based formulation of Effort is more relevant to laypeople’s fairness perception than traditional formulations based solely on the feature’s aggregate value (Romano, Bates, and Candes 2020; Ni et al. 2024). Finally, we demonstrate how to compute our EaF metric to audit AI models on two datasets, CLUE (criminal justice) and SHED (personal finance). [^2] Our supplementary materials (e.g., data, code) are uploaded to a Google Drive folder: https://drive.google.com/drive/folders/1RsSE1dUHDkn41iKCkIhY5QRdAVKVcrtB?usp=sharing
@@
In the technical CS/AI fairness literature, two main approaches mathematically formulate “fairness” based on the outcome of human/AI-assisted decision-making systems: group fairness and individual fairness. Group fairness, or group parity, is achieved when a statistical metric of interest (such as positive outcome rate or false positive rate) is equalized across different groups with respect to a sensitive feature, e.g., race or sex (Pedreshi, Ruggieri, and Turini 2008; Barocas, Hardt, and Narayanan 2019). Individual fairness, introduced formally by Dwork et al. (2012), is based on the intuition that similar individuals should get similar outcomes. Formally, for a pair of individuals, given an input-space distance metric to measure how differently situated they are and an output-space distance metric to measure how different the distributions of their possible outcomes are, individual fairness requires (as a constraint in an optimization problem) that their output-space distance should be upper-bounded by their input-space distance. The more similarly situated two persons are (e.g., same demographics and criminal histories), the more likely that they should receive similar outcomes (e.g., risk scores).  [^3] We discuss the legal relevance of group fairness and individual fairness under U.S. law in Appendix A.
@@
Recent AI fairness works attempt to incorporate some notion of “Effort” into fairness metrics by equalizing “Effort” across groups. They conceptualize “Effort” as the relationship between changes in mutable input features (e.g., prior arrests) and the probability predicted by a model of receiving a favorable outcome (e.g., granting bail). Approaches differ in terms of whether they take outcome parity as a requirement and attempt to equalize “Effort” across groups (Heidari, Nanda, and Gummadi 2019; Guldogan et al. 2023) or whether they take equalized “Effort” as a requirement and attempt to equalize outcome parity (Gupta et al. 2019; Huang et al. 2020; Von Kügelgen et al. 2022). These dual styles of approach are closely related to work on algorithmic recourse (Ustun, Spangher, and Liu 2019), in that they ask how much effort” an individual would have to put in to achieve a favorable outcome. Both approaches can be loosely considered to be in the space of counterfactual “Effort” and in contrast to our approach—we are interested in how much “Effort” an individual has already put in and use that as the value against which we seek parity.
Formally, existing approaches focus on counterfactual changes in input features ($\Delta\mathbf{x}$), which may lead to changes in the probability of receiving a favorable output label, e.g., from $f(\mathbf{x})<\gamma$ to $f(\mathbf{x}+\Delta\mathbf{x})\geq\gamma$ where $\gamma$ is a probability threshold. On the one hand, Heidari, Nanda, and Gummadi (2019) and Guldogan et al. (2023) set an upper threshold $\delta$ for the norm function $\mu$ of “Efforts”, i.e. $\mu(\Delta\mathbf{x})\leq\delta$, and seeks to equalize across groups the maximized favorable output rate. On the other hand, Gupta et al. (2019), Huang et al. (2020), and Von Kügelgen et al. (2022) set a lower threshold for the desirable output rate and seek to equalize across groups the minimized norm of “Efforts”. However, these works do not use actual changes over time, but they compare an individual with all others within the same group at only one time step to estimate $\Delta\mathbf{x}$ and $f(\mathbf{x}+\Delta\mathbf{x})$.  [^4] Shortcomings of the counterfactual Effort fairness literature and how our formulation overcomes them are in Appendix B.
@@
As reviewed by Massin (2017), the philosophy literature has used four main approaches to characterize Effort, of which two are directly related to Effort itself (rather than the perception of Effort): (1) Force-based account (how much force would it take to “move” an object; introduced by De Tracy (1803) and later formalized/advocated for by, i.a., Maine de Biran (2002)) and (2) Resource- or Energy-based account (spending Effort towards a goal means allocating part of a finite amount of resources, or energy tank, up until the energy tank is depleted; introduced by Arai (1912) and later formalized/advocated for by, i.a., Gendolla, Wright, and Richter (2012)).
The Force-based account addresses several problems unresolved by the Energy-based account, such as capturing the concept of “resistance” to Effort, distinguishing between “fatigue” and “Effort”, and explaining why Effort is unpleasant but praiseworthy. Massin (2017) informedly concludes that the Force-based account is “more fundamental”.
The three main Philosophy arguments by Massin (2017) in support of the force-based account are highly applicable to our AI fairness contexts: (1) As “only force-based accounts properly capture the idea of a resistance to our Effort,” these accounts can capture issues such as systemic discrimination that resists individual-level progress (Bohren, Hull, and Imas 2022). (2) As “force-based accounts, by contrast [to energy/resource-based accounts], stay clear of any substantive commitment about the relationship between fatigue and Effort,” it is compatible with our AI-assisted decision-making contexts where a decision subject with better Effort (e.g., fewer arrests or higher income after each year) does not necessarily suffer from more fatigue. (3) As “force-based accounts neatly explain the apparently paradoxical facts that we seem to praise and enjoy Efforts partly in virtue of their unpleasantness,” it is compatible with the later introduced inertia and acceleration terms in our force-based formulation of Effort because having grown up in childhood poverty but still managing to work harder for better income (or desistance) might be unpleasant but indicating of virtue and thus praiseworthy (McNeill and Farrall 2013). Bermúdez and Massin (2023) later expand this force-based conceptualization of Effort to conceptualize the “feeling of Effort”: “Among the possible Effort-first accounts, a force-based approach seems most promising, partly because it allows for a direct explanation of why the feeling of Effort involves the experience of countering a resistance.”
@@
To estimate “Effort” as an individual-level metric, we formulate the “more fundamental” Force-based account of “Effort” from philosophy (Massin 2017) by using Newton’s Second Law of Motion from Physics to develop an Effort-as-Force metric. Building on this idea, we formulate Effort-aware Fairness metrics to evaluate on real-world datasets.
@@
As Massin (2017) does not give specific mathematical formulae to compute Force or Effort, we turn to the Physics literature for the well-known Newton’s Second Law of Motion: $\mathbf{\mathbf{F}_{\text{net}}}=m\cdot\mathbf{a}$ where $\mathbf{F}_{\text{net}}$ is the net force applied on an object of inertia $m$ to give it an acceleration $\mathbf{a}$ (second-order time derivative of position $\mathbf{x}^{i}$) (Young, Freedman, and Ford 2014).555mm is more commonly known as “mass” in everyday Physics. Inertia is a more generalizable understanding of this metric. We look for analogies to model the two quantities $m$ and $\mathbf{a}$ to compute our Effort-as-Force metric $E\propto\mathbf{F}_{\text{net}}$. Assuming no friction, a person with inertia $m$ (characterized by systemic disadvantages or societal “holding-back” effects that are beyond their control) has to apply Effort $E$ to make their features (e.g., number of arrests or income) move in a desirable direction with an (observable) acceleration $\mathbf{a}$.
@@
Our first quantity to model is inertia $m$. We conceptualize $m$ as the societal “holding-back” effect and/or historical disadvantage against an individual beyond their control/agency. One example metric for inertia is hereditary disability, i.e., a disability passed down to a person by their parents. However, this feature might be hard to obtain (due to health privacy concerns) and difficult to quantify. Another example metric for inertia is the socio-economic background of a person’s family during their childhood, e.g., whether, back when the person was still a child, their family had to live below the federal poverty line, which is beyond the child’s control. In our criminal justice and personal finance contexts, childhood poverty was found to be associated with significantly higher “hazard rate of being convicted of violent criminality” (Sariaslan et al. 2014) and to have “quantitatively large detrimental effects” on adult earnings (Duncan, Ziol-Guest, and Kalil 2010).
@@
Next, acceleration $\mathbf{a}$ can be modelled as the second-order time derivative of task-relevant input features (e.g., number of past arrests in criminal justice, or income in personal finance). However, at first sight, there seems to be a problem with the second-order derivative. Suppose over a three-year period with equidistant time intervals ($\Delta t=\text{one year}=1$), a person’s annual income increases from $x_{0}=60k$ USD (first year) to $x_{1}=90k$ USD (second year) to $x_{2}=100k$ USD (third year). Our intuition suggests that they demonstrate positive Effort overall, as their annual income continued to increase, albeit at a slower velocity ($v_{0}=90k-60k=30k$ USD/year, followed by $v_{1}=10k$ USD/year). However, the slowing speed would result in a negative second-order derivative of $a_{0}=10k-30k=-20k$ (USD/year2), counterintuitively implying negative Efforts. Does this counterintuitive result mean the Force-based approach for Effort is wrong? Not necessarily. The underlying issue is that we should have modeled the temporal input feature (income) in a cumulative view: $X_{0}=60k$ USD, $X_{1}=60k+90k=150k$ USD, and $X_{2}=150k+100k=250k$ USD. In this cumulative formulation of income, velocity terms are $V_{0}=90k$ (USD/year) and $V_{1}=100k$ (USD/year); acceleration is $A_{0}=10k$ (USD/year2), intuitively implying positive Efforts. We use lowercase letters ($x,v,a$) to signal a non-cumulative view and uppercase letters ($X,V,A$) to signal a cumulative view. This example highlights the importance of transforming a non-cumulative feature into a cumulative feature before computing its acceleration to quantify Effort.
@@
We justify formally our choice of a cumulative formulation of the Effort-related feature to compute acceleration. A non-cumulative version of a feature can be obtained by differentiating (taking the derivative of) the cumulative version of that same feature. Therefore, acceleration (second-order derivative, i.e., derivative of the first-order derivative) in the cumulative version corresponds to velocity (first-order derivative) in the non-cumulative version, e.g., $A_{0}=v_{1}$. If we took the second-order derivative in the non-cumulative version, it would correspond to the third-order derivative in the cumulative version, which is theoretically at odds with force and hard to compute in practice.
Since Effort and AI-assisted decision making are social phenomena, we explain the social science relevance of our force-based formulation of Effort in Appendix C.
@@
To develop an Effort-aware Individual Fairness (EaIF) metric, we stipulate that similar individuals who spent similar Efforts should get similar outcomes. We propose a pairwise individual similarity function as a weighted combination between the aggregate task-relevant feature and Effort $E$. We use Definition 2.1 (Lipschitz mapping) by Dwork et al. (2012) as the individual fairness condition. Let $\mathbf{x}^{i}$ and $\mathbf{x}^{j}$ denote task-relevant feature vectors for a pair of individuals. A risk predictor $M$ maps individual $\mathbf{x}^{i}$ to a risk score, i.e., $M\mathbf{x}^{i}$. We define an Effort function $E(\cdot)$ that maps an individual $\mathbf{x}^{i}$ to an Effort score $E(\mathbf{x}^{i})$ and an aggregate (summing) function $S(\cdot)$ (e.g., to sum all non-cumulative values of income or past arrests $S(\mathbf{x}^{i})$ of individual $\mathbf{x}^{i}$). We also define an input-space individual distance function $d(\cdot,\cdot)$ over individuals’ feature representations and an output-space distance function $D(\cdot,\cdot)$ over their predicted risk scores. The predictor $M$ achieves perfect Effort-aware individual fairness if for every pair of ($\mathbf{x}^{i},\mathbf{x}^{j}$), this constraint is satisfied:
@@
$\quad\Longleftrightarrow\quad$ $\forall e_{0}\in\mathbb{R},\;\forall g_{i},g_{j}\in\mathcal{G},$
@@
where $e_{0},e_{0}+\epsilon]$ is a small range (bin) of Effort values, which can be computed from inertia $m$ and acceleration $\mathbf{A}$ (from the cumulative formulation of Effort-related feature). The terms $g_{i}$ and $g_{j}$ are demographic groups (e.g., White and Black), and $\mathcal{G}$ denotes the set of all such groups.
@@
For the following two empirical sections (human subjects experiment and metric demonstration pipelines), we use two real-world datasets that are relevant to two different AI-assisted decision-making contexts and have a temporal dimension. We also build simple predictive models to obtain AI predictions, on which Effort-aware Fairness can be evaluated.
@@
We select defendants from a real dataset, Client Legal Utility Engine (CLUE), with over $4.4$ million District Court and $700,000$ Circuit Court cases between $2012$ and $2020$ scraped by the Maryland Volunteer Lawyers Service (MVLS) from the Maryland Judiciary Case Search. [^6] https://casesearch.courts.state.md.us/casesearch/ CLUE contains criminal history and demographic features, e.g., race, sex, and age (at the time of arrest).
@@
For every defendant who had an arrest between $2016$ and $2018$, we develop simple machine learning models (Random Forest classifier, whose predictions we use for the human subjects experiment, and three other models, LightGBM, Logistic Regression, and Decision Tree, whose predictions we use for the metric demonstration) with 5-fold cross-validation to predict their (violent) recidivism risk (on a scale of 0 to $1$) within one year current their recent arrest. [^7] We provide more details about the predictive models in Appendix D.
@@
We use the “Survey of Household Economics and Decisionmaking” (SHED) published by the Federal Reserve since 2013.888www.federalreserve.gov/consumerscommunities/shed.htm To obtain the time dimension, we perform record linkage across four years ($2019$ to $2022$). After selecting only the households that have data available across all four years, we obtain the final set of $704$ households. Our Effort-related feature is the annual range of household income (I40). We turn this range-based feature into a number by taking the lower bound. Our outcome feature of interest is the respondent’s frequency of unpaid credit card balance (C4A), which can be of real-world predictive interest for financial institutions to process credit/loan applications.
@@
We must first decide on an appropriate population. Several major works in the AI fairness literature have used laypeople’s perception of fairness to decide on an appropriate fairness metric for a specific task, or even to construct new fairness metrics (Saxena et al. 2019; Harrison et al. 2020; Wang et al. 2019). We follow these precedents of informing our fairness metrics via laypeople’s fairness perception, based on two rationales. In principle, laypeople’s fairness judgment is vital for policy. Awad et al. (2020) argue that policymakers should be aware of and prepare for societal pushback of potentially beneficial but high-stakes decision-making AI tools. Liu, Du, and Li (2021) maintain that the discrepancy between a human-AI legal responsibility allocation framework from laws and the public’s expectations may hinder the adoption of and trust in AI at large. In practice, to potentially achieve statistical significance, a human subjects experiment typically needs hundreds of participants. For example, the mean sample size for online human subjects experiments at CHI, a leading Human-Computer Interaction conference, was 224 participants (Caine 2016). It is both expensive and logistically impractical to recruit such a large sample of technical CS/philosophy domain experts.